A soccer player kicks the baLL towards a goal that is 6.8 meters in front of him. The ball leaves his foot at 6 meters per second and an angle of 28 degrees above ground. Find the speed of the ball when the goalie catches it in front of the net.
1
Expert's answer
2012-10-09T09:15:53-0400
a=28
V0=6m/s
V=V0*cos(a)=6*0.89=5.34m/s